# Optimized Minimal 3D Gaussian Splatting

### v1

@@ Abstract
Figure 1: Our approach focuses on minimizing storage requirements while using only a minimal number of Gaussian primitives. By proposing an efficient attribute representation, including sub-vector quantization, we achieve scene representations under 5 MB with 600+ FPS rendering. We visualize qualitative examples (left) and the rate-distortion curve evaluated on the Mip-NeRF 360 dataset (right). All rendering speeds were measured on an NVIDIA RTX 3090 GPU, with values in parentheses in the left visualizations measured using an NVIDIA RTX 4090 GPU.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [26] has gained popularity for fast and photorealistic 3D scene reconstruction and rendering, offering a compelling alternative to conventional methods. By leveraging tile-based parallelism to approximate NeRF’s [37] volumetric rendering, 3DGS enables significantly accelerated rendering while maintaining high visual quality. This has facilitated a wide range of applications, such as dynamic scene reconstruction [58, 55], photorealistic avatar generation [45, 38], generative models [53, 10], and city-scale rendering [27, 46], demonstrating its versatility across various domains.
3DGS adjusts the number of Gaussian primitives during training by iteratively cloning or splitting Gaussians with high positional gradients while removing low-opacity Gaussians. However, this optimization process introduces a substantial number of redundant Gaussians (over 3 million per 360 scenes [2]), leading to excessive storage requirements and computational overhead. To address this issue, various approaches have been proposed, including pruning based on rendering loss [29, 33] or importance score [43, 12] and optimized densification strategies [36]. Notably, several methods [12, 13, 61] reduce the number of Gaussians to around 0.5 million, enabling real-time rendering even on low-capacity GPUs while preserving rendering quality.
Despite these efforts, reducing the number of Gaussians alone does not sufficiently mitigate storage overhead. Each Gaussian is parameterized by 59 learnable parameters, so even with a reduced number of primitives, storage consumption remains substantial (e.g., 133 MB in Fig. 1). To address this, many works have explored compressing Gaussian attributes by leveraging vector quantization [41, 11], neural fields [50], sorting mechanisms [39], and entropy optimization [7, 54], demonstrating considerable improvements in reducing storage consumption.
@@
In this paper, we propose Optimized Minimal Gaussian representation (OMG), an efficient compression framework that operates with a minimal number of primitives. To address the irregularity of sparse Gaussians and maximize the compressibility, we employ per-Gaussian features in a novel way. Although the reduced number of Gaussians leads to a decrease in local continuity, we can still leverage the spatial correlation associated with each Gaussian’s position. Therefore, we introduce a lightweight neural field model with negligible parameters to capture the coarse spatial feature. This feature is integrated with the per-Gaussian features to represent each attribute, as shown in Fig. 2. This approach requires fewer per-Gaussian parameters than directly learning the original attributes, enabling a more compact representation. While the proposed OMG architecture effectively represents sparse Gaussians, the use of per-Gaussian features impacts storage efficiency. To mitigate this, we introduce a sub-vector quantization (SVQ, Fig. 3(c)), which splits the input vector into multiple sub-vectors and applies vector quantization to each sub-vector. This approach alleviates the computational overhead associated with large vector quantization codebooks (Fig. 3(a)) and reduces the storage burden caused by the multiple indexing stages of residual vector quantization (Fig. 3(b)), while maintaining high-precision representation.
Finally, to retain only the minimal number of Gaussians, we introduce a novel importance metric that evaluates each Gaussian’s local distinctiveness relative to its neighbors, identifying the most informative Gaussians. This metric is used alongside existing importance scoring methods based on blending weights from training views [43, 12], further reducing the number of Gaussians while preserving scene fidelity.
Extensive experimental results demonstrate that OMG achieves a 49% reduction in storage compared to the previous state-of-the-art method [50], requiring only 4.1MB for the Mip-NeRF 360 dataset [2] while preserving comparable rendering quality. Additionally, OMG utilizes only 0.4 million Gaussians, enabling 600+ FPS rendering. These results underscore the effectiveness of OMG in both compression efficiency and computational performance, demonstrating it as a highly promising approach for 3D Gaussian Splatting representation.
@@ 2.1 Neural Radiance Fields
Neural Radiance Fields (NeRF) [37] introduced a pioneering approach for novel view synthesis by leveraging volumetric rendering in conjunction with Multilayer Perceptrons (MLPs) to model continuous 3D scenes. While NeRF achieves high-quality rendering, its reliance on MLP leads to inefficiencies, particularly in terms of slow training and inference times. To overcome these limitations, the following methods [14, 31] utilized explicit voxel-based representations, enabling significantly faster training compared to traditional MLP-based NeRF models. However, these approaches still suffer from slow inference speeds and impose substantial memory requirements, posing challenges for scalability and practical deployment in large-scale environments.
Compact representation. To mitigate the memory overhead while maintaining rendering fidelity, various works have been introduced, including grid factorization [5, 4, 15, 20, 16], hash grids [40, 6], grid quantization [52, 49], and pruning-based strategies [47]. Nevertheless, achieving real-time rendering for complex, large-scale scenes remains a formidable challenge. The fundamental limitation of these approaches stems from the necessity of dense volumetric sampling, which, despite optimizations, continues to constrain training and inference speed.
@@
Recently, 3D Gaussian Splatting (3DGS) [26] has emerged as a paradigm-shifting technique for real-time neural rendering by representing a scene with 3D Gaussian primitives. 3DGS leverages customized CUDA kernels and optimized algorithms to achieve unparalleled rendering speed while preserving high image quality. Unlike volumetric methods that require dense per-ray sampling, 3DGS projects Gaussians onto the image plane and rasterizes them tile-wise, significantly improving computational efficiency. Due to its versatility, 3DGS has become a dominant paradigm in 3D representation, leading to advancements across various domains and applications, such as mesh extraction [19, 22], simultaneous localization and mapping (SLAM) [25], dynamic scene representation [35], multi-resolution rendering [59], and further improvements in rendering quality [12]. However, 3DGS requires a substantial number of Gaussians to maintain high-quality rendering. Furthermore, each primitive is represented with multiple attributes, such as covariance matrices and spherical harmonics (SH) coefficients, requiring a large number of learnable parameters. Consequently, 3DGS demands substantial memory and storage resources, often exceeding 1GB per scene in high-resolution environments.
@@
Reducing the number of Primitives. To alleviate the substantial computational and memory overhead of 3DGS, numerous methods have been proposed to reduce the number of Gaussians while preserving rendering quality. Several approaches follow 3DGS by pruning low-opacity Gaussians, incorporated with opacity regularization [41], anchored Gaussians [34], or hyperparameter search [39]. An alternative approach utilized binary masking techniques [29, 7, 51, 54, 56], where pruning decisions are directly learned based on rendering loss. To optimize the binary masks, Compact-3DGS [41] initially adopted STE [3], while subsequent works [62, 33] employed Gumbel-Softmax.
@@
Another direction focuses on importance-based metrics to identify and remove redundant Gaussians. These methods primarily leverage each Gaussian’s blending weight contribution to rendering training-view images as a measure of importance [11, 12, 43, 17, 36]. LightGaussian [11] further incorporates Gaussian volume and opacity into the importance computation, while Taming 3DGS [36] integrates multiple information, including gradients, pixel saliency, and Gaussian attributes. Building upon these advancements, we introduce a novel importance metric that incorporates color distinction among neighboring Gaussians, enabling more effective selection of essential primitives.
@@
Attribute compression. Earlier methods employed conventional compression techniques such as scalar and vector quantization (VQ) [11, 41, 42, 39, 17, 44, 57] and entropy coding [42, 39, 29, 7, 8] to reduce storage requirements. VQ-based representations have proven highly efficient by the fact that many Gaussian attributes are redundant across a scene, allowing for compact encoding. However, a large codebook leads to substantial computational overhead, increasing training time. While residual vector quantization (R-VQ) [29] can alleviate computational costs, it introduces additional storage inefficiencies due to the need for multiple code indices.
@@
Another line of work explored structured representations, incorporating anchor-based encoding [34, 7, 56, 32] and factorization techniques [51], integrated with grid representations. Scaffold-GS [34] first introduced an anchor-based approach, where attributes of grouped Gaussians are encoded using shared anchor features and MLP-based refinements. Building upon this, subsequent methods [7, 9, 56, 32] incorporated context modeling to further improve compression rates. While showing high compression performance, these approaches require per-view processing, involving multiple MLP forward passes, which results in significant rendering latency.
@@
Recent efforts have utilized neural field architectures to exploit the local continuity of neighboring Gaussians. Compact-3DGS [29] encodes view-dependent color, while LocoGS [50] represents all Gaussian attributes except for view-independent color. However, unlike NeRF-based representations, where exact spatial positions are used as inputs to neural fields, mapping Gaussian center points to their corresponding attributes remains challenging. This difficulty leads to the use of large neural field models to achieve accurate reconstruction. In this work, we propose a novel approach that effectively captures both the continuity and irregularities across Gaussians, enabling a more efficient and compact attribute representation.
@@ 3.1 Overall Architecture
OMG is designed to accurately and efficiently represent the attributes of the minimal Gaussian primitives. Existing approaches [29, 50] have leveraged neural fields to exploit the local continuity of Gaussian attributes. However, as Gaussians become sparser, local continuity between them decreases. Unlike in dense representations, where smooth transitions between Gaussians can maintain fidelity, such transitions become less feasible in sparse settings. Especially for geometry, each Gaussian covers a larger spatial region, requiring a more specific scale and rotation to accurately capture structural details. Therefore, we retain the per-Gaussian parameterization for scale $s\in\mathbb{R}_{+}^{N\times 3}$ and rotation $r\in\mathbb{R}^{N\times 4}$ as in 3DGS.
@@
For appearance, local continuity can still be maintained even with increased sparsity. However, unlike NeRF, where the query input is a direct spatial point, mapping Gaussian center points to corresponding appearances is inherently challenging. This requires a larger neural field model to maintain high fidelity. Conversely, entirely disregarding local continuity leads to an inefficient representation, limiting the ability to capture meaningful spatial relationships. OMG addresses these challenges by integrating per-Gaussian attributes with a neural field structure, effectively leveraging both irregularity and continuity.
@@
To represent appearance, we learn the static feature $T\in\mathbb{R}^{N\times 3}$ and view-dependent feature $V\in\mathbb{R}^{N\times 3}$, as per-Gaussian attributes. As illustrated in Fig. 2, each feature is concatenated with the space feature $F_{n}$, derived from each Gaussian’s center position, to generate static and view-dependent color, and opacity. The space feature itself is efficiently parameterized using positional encoding and an MLP, ensuring a highly compact representation. Formally, this process can be expressed as follows:

### v2

@@ Abstract
Figure 1: Our approach focuses on minimizing storage requirements while using only a minimal number of Gaussian primitives. To achieve this, we introduce a local distinctiveness metric to identify the important Gaussians. Additionally, we propose a more efficient attribute representation, particularly for sparse Gaussians, by exploiting their irregularity and continuity. As a result, our method enables scene representation under 5 MB while achieving 600+ FPS rendering.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [3dgs] has gained popularity for fast and photorealistic 3D scene reconstruction and rendering, offering a compelling alternative to conventional methods. By leveraging tile-based parallelism to approximate NeRF’s [nerf] volumetric rendering, 3DGS enables significantly accelerated rendering while maintaining high visual quality. This has facilitated a wide range of applications, such as dynamic scene reconstruction [realtime4dgs, shapeofmotion], photorealistic avatar generation [3dgsavatar, human3dgs], generative models [dreamgaussian, text3dgs], and city-scale rendering [hierarch3dgs, octreegs], demonstrating its versatility across various domains.
3DGS adjusts the number of Gaussian primitives during training by iteratively cloning or splitting Gaussians with high positional gradients while removing low-opacity Gaussians. However, this optimization process introduces a substantial number of redundant Gaussians (over 3 million per 360 scenes [mip360]), leading to excessive storage requirements and computational overhead. To address this issue, various approaches have been proposed, including pruning based on rendering loss [c3dgs, maskgaussian] or importance score [radsplat, msv1] and optimized densification strategies [taming3dgs]. Notably, several methods [msv1, msv2, gaussianspa] reduce the number of Gaussians to around 0.5 million, enabling real-time rendering even on low-capacity GPUs while preserving rendering quality.
Despite these efforts, reducing the number of Gaussians alone does not sufficiently mitigate storage overhead. Each Gaussian is parameterized by 59 learnable parameters, so even with a reduced number of primitives, storage consumption remains substantial (e.g., 133 MB in Figure˜1). To address this, many works have explored compressing Gaussian attributes by leveraging vector quantization [compact3d, lightgaussian], neural fields [locogs], sorting mechanisms [c3d2d], and entropy optimization [hac, rdogaussians], demonstrating considerable improvements in reducing storage consumption.
@@
In this paper, we propose Optimized Minimal Gaussian representation (OMG), an efficient compression framework that operates with a minimal number of primitives. To address the irregularity of sparse Gaussians and maximize the compressibility, we employ per-Gaussian features in a novel way. Although the reduced number of Gaussians leads to a decrease in local continuity, we can still leverage the spatial correlation associated with each Gaussian’s position. Therefore, we introduce a lightweight neural field model with negligible parameters to capture the coarse spatial feature. This feature is integrated with the per-Gaussian features to represent each attribute, as shown in Figure˜2. This approach requires fewer per-Gaussian parameters than directly learning the original attributes, enabling a more compact representation. While the proposed OMG architecture effectively represents sparse Gaussians, the use of per-Gaussian features impacts storage efficiency. To mitigate this, we introduce a sub-vector quantization (SVQ, Figure˜3(c)), which splits the input vector into multiple sub-vectors and applies vector quantization to each sub-vector. This approach alleviates the computational overhead associated with large vector quantization codebooks (Figure˜3(a)) and reduces the storage burden caused by the multiple indexing stages of residual vector quantization (Figure˜3(b)), while maintaining high-precision representation.
Finally, to retain only the minimal number of Gaussians, we introduce a novel importance metric that evaluates each Gaussian’s local distinctiveness relative to its neighbors, identifying the most informative Gaussians. This metric is used alongside existing importance scoring methods based on blending weights from training views [radsplat, msv1], further reducing the number of Gaussians while preserving scene fidelity.
Extensive experimental results demonstrate that OMG achieves a 49% reduction in storage compared to the previous state-of-the-art method [locogs], requiring only 4.1MB for the Mip-NeRF 360 dataset [mip360] while preserving comparable rendering quality. Additionally, OMG utilizes only 0.4 million Gaussians, enabling 600+ FPS rendering. These results underscore the effectiveness of OMG in both compression efficiency and computational performance, demonstrating it as a highly promising approach for 3D Gaussian Splatting representation.
@@ 2.1 Neural Radiance Fields
Neural Radiance Fields (NeRF) [nerf] introduced a pioneering approach for novel view synthesis by leveraging volumetric rendering in conjunction with multi-layer perceptrons (MLPs) to model continuous 3D scenes. While NeRF achieves high-quality rendering, its reliance on MLP leads to inefficiencies, particularly in terms of slow training and inference times. To overcome these limitations, the following methods [fridovich2022plenoxels, nsvf] utilized explicit voxel-based representations, enabling significantly faster training compared to traditional MLP-based NeRF models. However, these approaches still suffer from slow inference speeds and impose substantial memory requirements, posing challenges for scalability and practical deployment in large-scale environments.
Compact representation. To mitigate the memory overhead while maintaining rendering fidelity, various works have been introduced, including grid factorization [tensorf, EG3D, kplanes, multensorf, strivec], hash grids [instant-ngp, howfar], grid quantization [VQAD, birf], and pruning-based strategies [masked]. Nevertheless, achieving real-time rendering for complex, large-scale scenes remains a formidable challenge. The fundamental limitation of these approaches stems from the necessity of dense volumetric sampling, which, despite optimizations, continues to constrain training and inference speed.
@@
Recently, 3D Gaussian Splatting (3DGS) [3dgs] has emerged as a paradigm-shifting technique for real-time neural rendering by representing a scene with 3D Gaussian primitives. 3DGS leverages customized CUDA kernels and optimized algorithms to achieve unparalleled rendering speed while preserving high image quality. Unlike volumetric methods that require dense per-ray sampling, 3DGS projects Gaussians onto the image plane and rasterizes them tile-wise, significantly improving computational efficiency. Due to its versatility, 3DGS has become a dominant paradigm in 3D representation, leading to advancements across various domains and applications, such as mesh extraction [sugar, 2dgs], simultaneous localization and mapping (SLAM) [splatam], dynamic scene representation [dynamic3dgs], multi-resolution rendering [mipsplatting], and further improvements in rendering quality [msv1]. However, 3DGS requires a substantial number of Gaussians to maintain high-quality rendering. Furthermore, each primitive is represented with multiple attributes, such as covariance matrices and spherical harmonics (SH) coefficients, requiring a large number of learnable parameters. Consequently, 3DGS demands substantial memory and storage resources, often exceeding 1GB per scene in high-resolution environments.
@@
Reducing the number of primitives. To alleviate the substantial computational and memory overhead of 3DGS, numerous methods have been proposed to reduce the number of Gaussians while preserving rendering quality. Several approaches follow 3DGS by pruning low-opacity Gaussians, incorporated with opacity regularization [compact3d], anchored Gaussians [scaffoldgs], or hyperparameter search [c3d2d]. An alternative approach utilized binary masking techniques [c3dgs, hac, f-3dgs, rdogaussians, contextgs], where pruning decisions are directly learned based on rendering loss. To optimize the binary masks, Compact-3DGS [compact3d] initially adopted STE [ste], while subsequent works [lp-3dgs, maskgaussian] employed Gumbel-Softmax.
@@
Another direction focuses on importance-based metrics to identify and remove redundant Gaussians. These methods primarily leverage each Gaussian’s blending weight contribution to rendering training-view images as a measure of importance [lightgaussian, msv1, radsplat, eagles, taming3dgs]. LightGaussian [lightgaussian] further incorporates Gaussian volume and opacity into the importance computation, while Taming 3DGS [taming3dgs] integrates multiple information, including gradients, pixel saliency, and Gaussian attributes. Building upon these advancements, we introduce a novel importance metric that incorporates color distinction among neighboring Gaussians, enabling more effective selection of essential primitives.
@@
Attribute compression. Earlier methods employed conventional compression techniques such as scalar and vector quantization (VQ) [lightgaussian, compact3d, morton3d, c3d2d, eagles, reduced3dgs, mesongs] and entropy coding [morton3d, c3d2d, c3dgs, hac, ff3d] to reduce storage requirements. VQ-based representations have proven highly efficient by the fact that many Gaussian attributes are redundant across a scene, allowing for compact encoding. However, a large codebook leads to substantial computational overhead, increasing training time. While residual vector quantization (R-VQ) [c3dgs] can alleviate computational costs, it introduces additional storage inefficiencies due to the need for multiple code indices.
@@
Another line of work explored structured representations, incorporating anchor-based encoding [scaffoldgs, hac, contextgs, compgs] and factorization techniques [f-3dgs], integrated with grid representations. Scaffold-GS [scaffoldgs] first introduced an anchor-based approach, where attributes of grouped Gaussians are encoded using shared anchor features and MLP-based refinements. Building upon this, subsequent methods [hac, hac++, contextgs, compgs] incorporated context modeling to further improve compression rates. While showing high compression performance, these approaches require per-view processing, involving multiple MLP forward passes, which results in significant rendering latency.
@@
Recent efforts have utilized neural field architectures to exploit the local continuity of neighboring Gaussians. Compact-3DGS [c3dgs] encodes view-dependent color, while LocoGS [locogs] represents all Gaussian attributes except for view-independent color. However, unlike NeRF-based representations, where exact spatial positions are used as inputs to neural fields, mapping Gaussian center points to their corresponding attributes remains challenging. This difficulty leads to the use of large neural field models to achieve accurate reconstruction. In this work, we propose a novel approach that effectively captures both the continuity and irregularities across Gaussians, enabling a more efficient and compact attribute representation.
@@ 3.1 Overall architecture
OMG is designed to accurately and efficiently represent the attributes of the minimal Gaussian primitives. Existing approaches [c3dgs, locogs] have leveraged neural fields to exploit the local continuity of Gaussian attributes. While effective in dense representations, this assumption weakens as Gaussians become sparser. In sparse settings, neighboring Gaussians are further apart, and smooth transitions between them become insufficient to maintain fidelity. Especially for geometry, each Gaussian covers a larger spatial region, requiring a more specific scale and rotation to accurately capture structural details. Therefore, we retain the per-Gaussian parameterization for scale $s\in\mathbb{R}_{+}^{N\times 3}$ and rotation $r\in\mathbb{R}^{N\times 4}$ as in 3DGS.
@@
For appearance, local continuity can still be maintained even with increased sparsity. However, unlike NeRF, where the query input is a direct spatial point, mapping Gaussian center points to corresponding appearances is inherently challenging. This requires a larger neural field model to maintain high fidelity. On the other hand, entirely disregarding local continuity leads to an inefficient representation, limiting the ability to capture meaningful spatial relationships. OMG addresses these challenges by integrating per-Gaussian attributes with a small neural field structure, effectively leveraging both irregularity and continuity.
@@
To represent appearance, we learn per-Gaussian attributes, namely a static feature $T\in\mathbb{R}^{N\times D}$ and a view-dependent feature $V\in\mathbb{R}^{N\times D}$, where $D$ is the dimensionality of each feature. As illustrated in Figure˜2, each feature is concatenated with the space feature $F_{n}$, derived from each Gaussian’s center position, to generate static and view-dependent color, and opacity. The space feature itself is efficiently parameterized using positional encoding and an MLP, ensuring a highly compact representation. Formally, this process can be expressed as follows:
$h_{n}^{(0)}=\text{MLP${}_{t}$}\big(\text{cat}(T_{n},F_{n})),\,o_{n}=\text{MLP${}_{o}$}\big(\text{cat}(T_{n},F_{n})),$ (3)
$h_{n}^{(1,2,3)}=\text{MLP${}_{v}$}\big(\text{cat}(V_{n},F_{n})\big),\,F_{n}=\text{MLP${}_{s}$}\big(\gamma(p_{n})\big),$ (4)
where $\text{cat}(\cdot,\cdot)$ denotes the concatenation function, $\gamma(\cdot)$ represents the positional encoding function, and $\text{MLP${}_{t}$}(\cdot),\text{MLP${}_{o}$}(\cdot),\text{MLP${}_{v}$}(\cdot),\text{MLP${}_{s}$}(\cdot)$ are the MLPs for static color, opacity, view-dependent color, and space feature, respectively.
3.2 Sub-vector quantization
Vector quantization (VQ) [vq] has shown high efficiency for representing Gaussian attributes, capitalizing on their inherently vectorized structure and strong global coherence across an entire scene. However, to maintain high fidelity, a large codebook size is required, inevitably resulting in substantial computational overhead and increased training complexity [soundstream] (Figure˜3(a)). To address these issues, Residual Vector Quantization (R-VQ) [soundstream] has been used as a hierarchical quantization strategy [c3dgs], progressively refining representations while reducing the size of each individual codebook. However, as shown in Figure˜3(b), multiple code indices per attribute result in increased storage overhead, illustrating a tradeoff between reducing per-codebook complexity and increasing overall storage requirements.
To navigate this tradeoff, we propose Sub-Vector Quantization (SVQ), which partitions the attribute vector into multiple sub-vectors and applies vector quantization separately to each component (Figure˜3(c)), motivated by Product Quantization [pq]. By reducing the dimensionality of each quantized unit, SVQ allows for smaller codebooks and more efficient lookups, which can balance computational cost and storage efficiency while maintaining high fidelity. We can apply SVQ to an input vector $z\in\mathbb{R}^{ML}$, where $M$ and $L$ represent the total number of sub-vectors (partitions) and the sub-vector length, respectively. Each partition $m\in\{1,...,M\}$ has an independent codebook $C^{(m)}\in\mathbb{R}^{B\times L}$, where $B$ denotes the number of codewords per codebook. The codeword selection is based on the nearest match from $C^{(m)}$, with $C^{(m)}[j]$ representing the $j$-th codeword corresponding to the $m$-th sub-vector. More formally, SVQ-applied vector $\hat{z}$ can be formulated as follows,
$i_{m}=\arg\min_{j}\|z_{m}-C^{(m)}[j]\|^{2}_{2},\quad m\in\{1,...,M\},$ (6)
SVQ ensures significantly reduced computation with small codebooks compared to VQ. We apply SVQ to geometric attributes $s_{n},r_{n}$, resulting in quantized vectors $\hat{s}_{n},\hat{r}_{n}$, which are then used for 3DGS rendering. For the appearance features $T_{n},V_{n}$, we first concatenate them and apply SVQ. The resulting quantized features are then split back into two components $\hat{T}_{n},\hat{V}_{n}$, which replace $T_{n}$ and $V_{n}$ in Equations˜3 and 4.
3.3 Local distinctiveness for important scoring
OMG adopts importance scoring to identify essential Gaussians and retain a minimal number of them. Existing scoring-based pruning methods typically determine the importance of each Gaussian based on its blending weights (the values multiplied by $c_{k}$ in Equation˜1) across training-view renderings. We use two factors as our baseline metric: (1) whether it has been the most dominant contributor for at least one ray [msv1, msv2] and (2) its total blending weight contribution across all training rays [radsplat, eagles]. Formally, we define the base importance score $\bar{I}$ as:
To mitigate these issues, we propose incorporating a local distinctiveness metric into the importance computation. Specifically, we introduce an additional term that measures the similarity of the static appearance feature $T\in\mathbb{R}^{N\times D}$ between neighboring Gaussians, ensuring that locally distinct Gaussians show high importance. The final importance score is defined as:
$I_{i}=\bar{I}_{i}\left(\frac{1}{D}\sum_{j\in\mathcal{N}_{i}^{K}}\|T_{i}-T_{j}\|_{1}\right)^{\lambda},$ (8)
where $\mathcal{N}_{i}^{K}$ denotes the set of $K$-nearest neighbors of Gaussian $i$, and $\lambda$ is a scaling factor that adjusts the sensitivity to appearance variation. As computing exact $K$-nearest neighbors for every Gaussian is computationally expensive, we approximate neighbor selection by sorting Gaussians in Morton order and selecting Gaussians with adjacent indices as their local neighbors. We remove low-importance Gaussians using CDF-based thresholding [compress1mb] with a threshold $\tau$.
Figure 4: Qualitative results of OMG compared to 3DGS and LocoGS with COLMAP initialization. We provide per-image rendering PSNR with storage requirements for each scene.
4 Experiment
4.1 Implementation details
Following the previous works, we evaluated our approach on three real-world datasets, Mip-NeRF 360 [mip360], Tanks&Temples [tnt], and Deep Blending [db]. Our model is implemented upon Mini-Splatting [msv1], one of the methods achieving high performance with a small number of Gaussians. We have conducted simple post-processings after training: 1) Applying 16-bit quantization to the position and compressing with G-PCC [gpcc]. 2) Huffman encoding [huffman] to SVQ indices. 3) Storing all the components into a single file with LZMA [lzma] compression. We provide five OMG variants (XS, X, M, L, XL), adjusting storage requirements. The only factor controlling the storage is the CDF-based threshold value $\tau$ of Gaussian importance, which is set to 0.96, 0.98, 0.99, 0.999, and 0.9999 for each variant, respectively. Further implementation details are provided in the supplementary materials.
Table 1: Quantitative results of OMG evaluated on the Mip-NeRF 360 dataset. Baseline results are sourced from the LocoGS [locogs] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU. LocoGS* refers to LocoGS initialized with COLMAP, instead of dense Nerfacto initialization. We highlight the results among compression methods by coloring the best, second-best, and third-best performances.
Mip-NeRF 360
3DGS 27.44 0.813 0.218 822.6 127
Scaffold-GS [scaffoldgs] 27.66 0.812 0.223 187.3 122
Mini-Splatting [msv1] 27.39 0.822 0.216 119.5 (601)
CompGS [compact3d] 27.04 0.804 0.243 22.93 236
Compact-3DGS [c3dgs] 26.95 0.797 0.244 26.31 143
C3DGS [morton3d] 27.09 0.802 0.237 29.98 134
LightGaussian [lightgaussian] 26.90 0.800 0.240 53.96 244
EAGLES [eagles] 27.10 0.807 0.234 59.49 155
SOG [c3d2d] 27.01 0.800 0.226 43.77 134